The composer Beethoven wrote 9 symp S plano concertos (music for plano and orchestra), and 32 plano sonatas (music for solo piano) (a) How many ways are there to play first a Beethoven symphony and then a Beethoven piano concerto ways (b) The manager of a radio station decides that on each successive evening (7 days per week), a Beethoven plano concerto will be played followed by a Beethoven piano sonata followed by a Beethoven symphony. For how many years could this policy be continued before exactly the same program would have to be repeated? (Assume there are 365 days in a year. Round your answer up to the nearest whole number,) years
Answer #1

SOlution:-
(a) There are 45 ways.

=> 9*5 = 45

(b) n = (9*5*32) = 1440 days

Number of years = 1440/365 = 3.94 (1 year = 365 days)

= 4 years.